OpenStudy (anonymous):

can some one help me solve 2x(x-3)=3x+2. this is a completeing the square problem. I have posted it 3 times and I have gotten 3 different answers. Im not sure which one is correct. please help?

OpenStudy (anonymous):

4x^2-6x-3x=2 4x^2-9x-2=0

OpenStudy (anonymous):

9+7 / 8 or 9-7 /8

OpenStudy (anonymous):

x=2 ** x=1/4

OpenStudy (anonymous):

see this answer is different from the other 3. So would you be able to work the entire problem? show work?